OpenStudy (anonymous):

Fred saved two thirds as much money as John. John saved 5 times as much as Mary. Together they saved $14. How much did each save? please helpp..

OpenStudy (anonymous):

please help cause im not suree :/

OpenStudy (anonymous):

Can you formulate an equation?

OpenStudy (anonymous):

no i need help on that. :/

OpenStudy (apoorvk):

Okay let's work this out! So, let the amount saved by Mary be '\(x\)' dollars. So, since John saved 5 times as much as Mary did, he saved times '5 times x' = '\(5x\)' dollars. And. Fred saved two-thirds of what John saved, so 'two-thirds of 5x' = \(\frac 2 3 \times 5x\) = \(\frac{10x} 3\) Now as the questions states, the total amount saved was 14 dollars. So, \(\large x + 5x + \frac{10x}3 = 14\) Can you solve this and find 'x'?

OpenStudy (anonymous):

so what is Fred's number?

OpenStudy (anonymous):

is it just 10x/3?

OpenStudy (apoorvk):

You know Fred's savings is 10x/3. So to find out what it exactly is, you must know the value of 'x', right?

OpenStudy (anonymous):

yeah now i noticed

OpenStudy (anonymous):

i just don't know if i should change the fraction into a decimal

OpenStudy (apoorvk):

And you have simple equation formulated above, using which you can find out 'x'. The equation: x+5x+(10x/3)=14 You need to add the terms in 'x', and then solve it. No just let it remain as a fraction, that would help you in the calculations. Take the LCM and ad them up - can you do that?

OpenStudy (anonymous):

um i don't know :(

OpenStudy (anonymous):

i don't know what to do when there is a fration

OpenStudy (apoorvk):

Well, if you know how to add it in decimal system, then use that method itself - no penalties. :]

OpenStudy (apoorvk):

10x/3 = 3.33 x ---> do you seem to be getting this?

OpenStudy (anonymous):

yeah to me decimals make it easier than a fraction

OpenStudy (anonymous):

but i still can't get the answer :/

OpenStudy (apoorvk):

No worries... so, you do understand that 10x/3 = 3.33x right?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

do you combine x and 5x?

OpenStudy (apoorvk):

Okay, so we have - x + 5x + 3.33x = 14 So, what's (x + 5x + 3.33x) ??

OpenStudy (apoorvk):

Yeah exactly! I *combine* the like terms - that is x and 5x and 3.33x!

OpenStudy (anonymous):

wait then do you subtract?

OpenStudy (anonymous):

cause when i do that i don't want to get a decimal as an answer

OpenStudy (anonymous):

it feels like it doesn't make sense

OpenStudy (apoorvk):

Nope, why would I subtract!! And no you have no choice, you would get a decimal as an answer. :| But don;t worry it isn't half as difficult as it looks ;] x + 5x + 3.33x taking 'x' common, =x(1 + 5 + 3.33) Can you add 1,5, and 3.33??

OpenStudy (anonymous):

its 4.83

OpenStudy (anonymous):

helloo?

OpenStudy (apoorvk):

Nope. 1 + 5 + 3.33 ---> use a calculator.

OpenStudy (anonymous):

i know its 9.33

OpenStudy (anonymous):

and the you diivide it by 14

OpenStudy (anonymous):

i get it and is the answer 0.6?

OpenStudy (anonymous):

and then you substitute the others

OpenStudy (apoorvk):

Exactly! Now you get it!!! Wonderful!

OpenStudy (anonymous):

ok thank you! finally! :D

OpenStudy (apoorvk):

Just that 9.33/14 = 0.67 more exactly.

OpenStudy (apoorvk):

You're welcome! :]

OpenStudy (anonymous):

thank you so much!